Status Report

NOAA GOES SATOPS Morning Report 19 Dec 2000

By SpaceRef Editor
December 19, 2000
Filed under ,

SATOPS MORNING REPORT – December 19, 2000

POES

No change in the status of the POES spacecraft.

NOAA-14  rev 30774 / W  at 0854Z on December 19:  G1B
was noisy and 96 % was recovered.  Recorder management sequence constraints
prevent further attempts at recovery.

NOAA-15 rev 13510 / W at 1402Z on December 18:  HIRS filter wheel
motor commanded to low power.

NOAA-15 rev 13516 & 13517 / W at 2341Z on December 18:  Commanded
spacecraft to switch Search and Rescue Repeater (SARR) receiver to side
B.

POES operations were nominal over the past 24 hours.

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

GOES operations were nominal over the past 24 hours.

 

Questions or comments to [email protected]

SpaceRef staff editor.